Q.1 Using the table below, answer the questions that follow it.

A

B

C

D

E

Pizza

0

1

2

3

4

Robots

20

18

14

8

0

  1. What are the assumptions for a given production possibilities curve?
  2. What is the opportunity cost of one pizza when moving from B to point C? What moving from D to point E?
  3. Do these combinations demonstrate constant or increasing opportunity costs?
Answer #1

a. The assumptions for a given production possibilities curve:

For a given period, the different combinations are drawn assuming that the resources, technology and production technique in the economy are held constant.

b. The opportunity cost of one pizza, when moving from point B to point C, is 4 units of robots.

18 robots – 14 robots = 4 robots

The opportunity cost of one pizza, when moving from point D to point E, is 8 units of robots.

8 robots – 0 robots = 8 robots

c. These combinations demonstrate increasing opportunity costs, as more is produced. (observe the last column of the table, MOC is increasing)

Combinations

Pizza

Robots

Marginal opportunity cost

A

0

20

---

B

1

18

2/1=2

C

2

14

4/1=4

D

3

8

6/1=6

E

4

0

8/1=8

Marginal opportunity cost = No. of units sacrificed / No. of units gained